Lt 1/x--> 0 {cos√(x+1) - cos√x}Pls answer as soon as possible

Grade:12th pass

1 Answers

jagdish singh singh
7 years ago
\hspace{-0.7 cm}l=\lim_{x\rightarrow \infty}\cos \sqrt{x+1}-\cos \sqrt{x} = \lim_{x\rightarrow \infty}\left(\frac{\cos \sqrt{x+1}-\cos \sqrt{x}}{x+1-x}\right)$\\\\\\ So we get $l= (\cos \sqrt{x})` = \lim_{x\rightarrow \infty}\frac{\sin \sqrt{x}}{\sqrt{x}} = 0$